OpenStudy (anonymous):

i need help with this problem 4(4)-2(5)

OpenStudy (anonymous):

4x4 = 16 2x5=10 16-10 = 6

OpenStudy (toxicsugar22):

16-10=6

OpenStudy (toxicsugar22):

you do 4*4= the 2*5=

OpenStudy (anonymous):

Parentheses means multiply when it's right next to something, so 4*4-2*5=16-10=6.